When 163, 216 and 587 are in turn divided by certain number, the remainder is 4 in each case. Find the number
Answer by Alan3354(69443)

When 163, 216 and 587 are in turn divided by certain number, the remainder is 4 in each case. Find the number
-----------
163 - 4 = 159
216 - 4 = 212
587 - 4 = 583
--------------
Find the common factor